# Towards Robust Assessment of Pathological Voices via Combined Low-Level Descriptors and Foundation Model Representations

### v3

@@ 4.1 Dataset
The audio samples used in this study come from the Perceptual Voice Quality Database (PVQD) provided by The Voice Foundation [4]. This dataset contains 296 audio recordings, each featuring one person’s sustained /a/ and /i/ vowels and continuous speech sentences defined by the CAPE-V protocol, providing a mixture of controlled and naturalistic speech material. The audio files are stored in WAV format, encoded with 16-bit resolution and a sampling rate of 44.1 kHz, ensuring high-fidelity acoustic analysis. In addition to raw audio, the dataset provides rich metadata including the speaker’s age, gender, diagnosis, and expert perceptual ratings using the CAPE-V and GRBAS frameworks. CAPE-V scores are continuous, ranging from 0 to 100, and are used to assess dimensions such as overall severity, breathiness, and strain. GRBAS scores use a 4-point ordinal scale (0–3) to describe Grade, Roughness, Breathiness, Asthenia, and Strain  [33]. Each recording was independently rated by two qualified speech-language pathologists, and the final perceptual score for each dimension was the average of the two raters to ensure reliability and reduce subjective bias. Table 1 summarizes the score distribution of each dimension of the two scales.
Table 1: Descriptive Statistics of CAPE-V and GRBAS Ratings in the PVQD Dataset.
Scale Attribute Mean Median Mode Min Max
CAPE-V (0–100) Severity 29.4 19.5 19.3 0.33 98.67
Roughness 20.7 13.7 9.7 0.17 84.83
Breathiness 19.8 12.2 5.0 0.00 99.50
Strain 21.1 12.2 4.5 0.12 96.83
Pitch 16.3 9.3 0.5 0.00 99.17
Loudness 18.7 8.8 0.7 0.00 99.17
GRBAS (0–3) Grade 1.0 0.8 0 0 3
Roughness 0.8 0.7 0 0 3
Breathiness 0.7 0.4 0 0 3
Asthenia 0.6 0.2 0 0 3
Strain 0.8 0.5 0 0 3
@@ 4.2 Data Split for Training and Testing
To prevent data leakage and ensure generalizability, the PVQD dataset was split at the patient level, meaning that each speaker was exclusively assigned to either the training set or the test set. This ensured that the model was evaluating pathological voice quality on unseen speakers rather than memorized speech patterns, thus validating its ability to generalize beyond the training set. This approach follows that used in [17] to ensure consistency with previous studies. The PVQD dataset originally contained 296 recordings, but 13 corrupted files were excluded, leaving a total of 283 valid samples. Specifically, 226 samples were used for training and 57 samples for testing. Table 2 provides the demographic distribution of speakers in the training and test sets, including gender ratio and age range. For a more comprehensive evaluation, vowel segments and continuous speech segments were extracted from each recording. Therefore, the PVQD dataset was divided into two subsets: PVQD-A (vowel-only subset), which contains /a/ vowel segments; and PVQD-S (speech-based subset), which consists of continuous speech segments. In this way, vowel phonation (PVQD-A) and continuous speech (PVQD-S) were evaluated independently. The number of samples in each subset is shown in Table 3. Since each recording contains one /a/ vowel and multiple continuous speech segments, there are more samples of continuous speech than /a/ vowel.
@@ 4.6 Noise Robustness Setup
We introduced a noise-augmented version of the PVQD dataset to evaluate the robustness of the model under adverse acoustic conditions. Table 4 summarizes the noise types and signal-to-noise ratio (SNR) levels used for training, seen test scenarios, and unseen test scenarios. Each set contains the original clean utterances. The training set was augmented with four noise types: white noise, pink noise (colored), cafeteria babble, and cocktail party (background), with SNRs of -5 dB, 0 dB, 5 dB, and 10 dB, respectively. This combination includes both stationary and non-stationary interference commonly used in speech robustness studies. To evaluate generalization, we introduced unseen noise types into the evaluation process: brown noise (low-frequency), baby cry, and laughter, the latter two representing non-speech vocalization of emotional expression. These noise types were chosen to reflect real-world clinical and telehealth environments, where voice assessment may be performed in varied background settings. Fig. 2 illustrates the waveforms and spectrograms of the PVQD-A and PVQD-S subsets under different noise types. The selected SNR levels (–5 dB to 10 dB) cover a wide range of environments from challenging to moderate noise, aligning with previous research in the field of ASR and speech enhancement, ensuring that the robustness evaluation reflects the real-world clinical environment.
@@ 5.1 Comparison of VOQANet with the Baseline
Table 5 shows the utterance-level evaluation results of VOQANet with different features (including the embeddings of the last layer of WavLM or Whisper) on CAPE-V and GRBAS prediction. We compare VOQANet with the methods of Lin et al. [17], which include traditional or neural regressors, using LLDs, traditional signal transformations (mel-frequency cepstral coefficient and modulation spectrum (MFCC+MS)), raw waveform, and features extracted by pre-trained models (e.g., Wav2Vec2 (W2V2), HuBERT, WavLM, and Whisper). Lin et al. only evaluated their methods on CAPE-V prediction on the PVQD-A subset. Their results show that traditional features (including HF and MFCC+MS) outperform SFM features. SFM features performed worse than expected, except for the representation of the last layer of Whisper (Whisper (Last)). The lowest RMSE is 14.76 and the highest PCC is 0.69. However, their best results are significantly worse than those of VOQANet, suggesting that these traditional features are not sufficient to model the complex acoustic patterns relevant to pathological voice assessment. Moreover, their models were trained using data augmentation with noise injection, whereas our VOQANet models in this experiment was trained on the original PVQD-A dataset without data augmentation. Despite this difference, VOQANet still performs significantly better than these baselines. On PVQD-A, VOQANet reduces RMSE to 10.514 (Whisper (Last)) and 9.955 (WavLM (WS)), and the latter is more than 30% lower than the best baseline. PCC is also improved substantially, exceeding 0.8 for both embeddings.
VOQANet with WavLM (Last) outperforms VOQANet with Whisper (Last) on CAPE-V prediction on the PVDQ-S subset, suggesting that WavLM is particularly well suited to capture prosodic and phonetic nuances in continuous speech. Similar improvements were observed for GRBAS prediction on both sunsets. While the numerical gains in GRBAS prediction may appear more subtle in absolute terms compared to CAPE-V prediction due to the narrower range of GRBAS (0-3), these gains are meaningful from a clinical perspective.
@@ 5.2 Comparison of SFM Types and Representations
In this study, WavLM and Whisper are used as backbone SFMs because of their excellent performance in speech quality assessment tasks [31]. Another study also highlighted the growing role of self-supervised learning models in non-intrusive speech assessment [38]. To further analyze the effectiveness of SFM-based representations, we compare two adaptation strategies: last-layer features (Last) and weighted-sum aggregation (WS) in Whisper and WavLM. This experiment adopts utterance-level evaluation. As shown in Table 6, the weighted-sum representation consistently outperforms the last-layer representation in all configurations. For example, for GRBAS prediction on the PVQD-A subset, WavLM WS yields lower RMSE (0.369 vs. 0.380) and higher PCC (0.809 vs. 0.795) than its last-layer counterpart. This is consistent with previous findings in the speech processing literature [18], [19], which show that intermediate transformer layers encode diverse and complementary representations, some of which better capture prosodic and phonatory features critical to the perception of pathological voice quality.
@@
From Table 6, we can also see that WavLM outperforms Whisper in CAPE-V and GRBAS predictions on both the PVQD-A and PVQD-S subsets. In particular, on the PVQD-S subset, WavLM achieves lower RMSE (9.209 vs. 9.933) and higher PCC (0.870 vs. 0.863) than Whisper when using WS features for CAPE-V prediction. The performance gap is even more pronounced in GRBAS prediction, with WavLM WS achieving a PCC of 0.845 on PVQD-S, compared to 0.822 for Whisper WS. These findings confirm that WavLM’s denoising pretraining and fine-grained acoustic modeling have advantages in disordered voice settings, especially when handling longer or more naturalistic speech. Overall, this experiment confirms that WavLM (WS) provides the most informative and powerful representation and will be used in subsequent experiments.
5.3 Comparison between VOQANet and VOQANet+
While SFM (e.g., WavLM) embeddings are effective in capturing high-level acoustic and prosodic information, they are not explicitly optimized for clinically salient voice quality traits. To address this limitation, VOQANet+ incorporates LLDs (jitter, shimmer, and HNR, or JSH for short). These features have been shown to correlate with pathological voice quality dimensions and are widely adopted in clinical voice analysis and speech processing systems due to their ability to reflect phonatory stability and noise characteristics [39]. Combined with WavLM embeddings, these LLDs provide interpretable low-level signal-based information that complements deep learning-based representations.
As shown in Table 7, VOQANet+ consistently achieves the lowest RMSE and highest PCC across all tasks and evaluation levels, outperforming VOQANet with SFM alone. The performance improvement is particularly prominent at the patient level, where predictions are aggregated over utterances from the same speaker to mimic real-world assessment. For example, in GRBAS prediction on PVQD-S, VOQANet+ improves the patient-level PCC from 0.867 to 0.874 and reduces RMSE from 0.297 to 0.289. These results indicate that adding LLDs helps to supplement low-level acoustic cues, such as irregularities in frequency or amplitude, that may not be adequately captured by SFM embeddings alone.
Fig. 3 visualizes the CAPE-V scores predicted by VOQANet versus the actual scores, while Fig. 4 does so for GRBAS. Fig. 5 visualizes the CAPE-V scores predicted by VOQANet+ versus the actual scores, while Fig. 6 does so for GRBAS. In each figure, the x-axis denotes the ground-truth scores assigned by expert raters, and the y-axis represents the model’s predicted scores. The top row (e.g., Fig. 3(a–b)) corresponds to utterance-level predictions, and the bottom row (e.g., Fig. 3(c–d)) shows patient-level predictions obtained by averaging the prediction scores of different utterances of the same speaker. Each point represents an utterance (or speaker), the red line shows the regression fit, and the shaded area is the 95% confidence interval. VOQANet+ shows tighter clustering near the diagonal, especially in the patient-level plots (e.g., Fig. 3(d) vs. Fig. 5(d) and Fig. 4(d) vs. Fig. 6(d)), which indicates stronger prediction consistency and lower variance. These findings further demonstrate that LLDs can enhance the model’s ability to estimate perceptual ratings, especially in cases of higher severity, where expert judgments tend to be more variable. In summary, by combining black-box deep learning-based representations with interpretable acoustic features, VOQANet+ improves the clinical relevance, trustworthiness, and robustness of pathological voice quality prediction.
5.4 Utterance-Level Evaluation vs. Patient-Level Evaluation
In clinical settings, auditory perceptual judgments are often made by listening to multiple utterances from a single speaker. To reflect this practice, we evaluate model performance at both the utterance level and the patient level by averaging the predictions for all utterances from the same speaker. As shown in Table 7 and visualized in scatter plots (Figs. 3, 4, 5, 6), patient-level evaluation achieves higher PCC and lower RMSE than utterance-level evaluation. For example, when using VOQANet+ (with WavLM (WS) + JSH features) for CAPE-V prediction on PVQD-S, PCC improved from 0.883 at the utterance level to 0.908 at the patient level, and RMSE decreased from 8.720 to 7.356.
To further investigate the contribution of different speech types, we compare the results of VOQANet+ (with WavLM (WS) + JSH features) on the PVQD-A (vowel-based) and PVQD-S (sentence-based) subsets. As shown in Fig. 7, sentence-based predictions on PVQD-S consistently outperform vowel-based predictions on PVQD-A, especially at the patient level. This supports the hypothesis that longer continuous speech utterances provide richer prosodic and articulatory information, allowing the model to make more accurate and stable predictions. These results suggest that while sustained vowels are still clinically useful, sentence-level inputs can provide more contextual acoustic dynamics, such as stress, pitch variation, and connected phonation, which are highly informative for complex perceptual dimensions such as strain or roughness.
Moreover, the advantage of sentence-based predictions is further amplified when LLDs are incorporated into the model (see the comparison of VOQANet and VOQANet+ in Table 7). This suggests a synergistic effect between rich SFM-derived representations and domain-informed acoustic markers, especially in the capture of the characteristics of voice disorders. These findings also have important implications for robustness. As we will explore in the next subsection, sentence-level input enables VOQANet+ to maintain stronger performance under seen and unseen noisy conditions. This resilience further highlights the clinical value of incorporating continuous speech into the automated voice assessment framework.
5.5 Robustness to Seen and Unseen Noise
To examine the generalizability of the models under adverse acoustic conditions, we tested the robustness of VOQANet and VOQANet+ under various noisy conditions on the noise-augmented PVQD dataset. To improve the robustness of dysphonic voice detection, recent approaches have focused on generating acoustic feature embeddings that are sensitive to vocal quality and robust across different corpora [40]. This emphasis on real-world acoustic variability and noise resilience stems from recent research in voice disorder modeling, which highlights the importance of domain robustness for clinical applicability [41]. As shown in Table 8, both models were evaluated for seen (i.e., white, pink, babble, and cocktail party noise used during training) and unseen (i.e., baby cry, laughter, and brown noise not included in training) noise types at multiple SNR levels from -5 dB to 10 dB. On both CAPE-V prediction and GRBAS prediction tasks, VOQANet+ consistently outperforms VOQANet under noisy conditions, demonstrating greater resilience to both seen and unseen disturbances. For example, for patient-level GRBAS prediction on PVQD-S under unseen noise, VOQANet+ achieves a PCC of 0.855, compared to 0.832 for VOQANet, while reducing RMSE from 0.336 to 0.313. Similarly, for utterance-level CAPE-V prediction on PVQD-S, VOQANet+ improves PCC from 0.809 to 0.811 and reduces RMSE from 10.627 to 10.579. While numerically modest, these improvements are meaningful in noisy clinical or telehealth scenarios, where the reliability of scoring under real-world background conditions is critical.
This improvement highlights that LLD (JSH), rooted in perturbation measures and periodicity detection, provides complementary information that is robust to noise and remains stable even when deep SFM embeddings degrade. By capturing signal-level voice irregularities that are less susceptible to spectral masking or background interference, VOQANet+ benefits from richer and more noise-resistant representations. Moreover, VOQANet+ shows less performance degradation from seen to unseen noise than VOQANet, especially in GRBAS prediction. This highlights the advantage of combining interpretable clinically meaningful features with learned embeddings to form a more adaptable model that maintains prediction quality even when faces with out-of-distribution acoustic environments. The results confirm that VOQANet+ provides a more robust solution for deployment in noisy real-world environments.

### v4

@@ 4.1 Dataset
The audio samples used in this study come from the Perceptual Voice Quality Database (PVQD) provided by The Voice Foundation [4]. The dataset includes 296 recordings, each containing sustained /a/ and /i/ vowels and connected-speech samples following the six standardized CAPE-V sentences [7]. These sentences are part of the official CAPE-V protocol and do not involve number counting or spontaneous speech, designed to elicit diverse phonetic contexts and form part of the official CAPE-V protocol, ensuring the material reflects clinically relevant speech patterns. Combining sustained vowels with CAPE-V sentences aligns with standard dysphonia assessment practices, capturing both glottal-source characteristics and connected-speech dynamics. All recordings are stored in 16-bit WAV format at a 44.1 kHz sampling rate. In addition to raw audio, the dataset also provides metadata on age, gender, diagnosis, and expert perceptual ratings for both CAPE-V and GRBAS. CAPE-V scores are continuous, ranging from 0 to 100, and are used to assess dimensions such as overall severity, breathiness, and strain. Although CAPE-V has been revised to CAPE-Vr, the PVQD corpus provides annotations based on the original CAPE-V framework, so these CAPE-V annotations were adopted in this study. GRBAS scores use a 0–3 ordinal scale [40] and are provided for both vowel (PVQD-A) and connected-speech (PVQD-S) samples, enabling direct comparison across phonation tasks. This represents a data-driven extension of the GRBAS framework beyond its traditional clinical use. Each recording was independently rated by two qualified speech-language pathologists, and the final perceptual score for each dimension is the average of the two raters to ensure reliability and reduce subjective bias. Table 1 summarizes the score distribution of each dimension of the two scales.
Table 1: Descriptive Statistics of CAPE-V and GRBAS Ratings in the PVQD Dataset (N=296)
. Scale Attribute Mean Median Mode Min Max CAPE-V (0–100) Severity 29.4 19.5 19.3 0.33 98.67 Roughness 20.7 13.7 9.7 0.17 84.83 Breathiness 19.8 12.2 5.0 0.00 99.50 Strain 21.1 12.2 4.5 0.12 96.83 Pitch 16.3 9.3 0.5 0.00 99.17 Loudness 18.7 8.8 0.7 0.00 99.17 GRBAS (0–3) Grade 1.0 0.8 0 0 3 Roughness 0.8 0.7 0 0 3 Breathiness 0.7 0.4 0 0 3 Asthenia 0.6 0.2 0 0 3 Strain 0.8 0.5 0 0 3
  • Note. Each attribute was rated across all available recordings (N=296)
@@ 4.2 Data Split for Training and Testing
To prevent data leakage and ensure generalizability, the PVQD dataset was split at the patient level, meaning that each speaker was exclusively assigned to either the training set or the test set. This ensured that the model was evaluating pathological voice quality on unseen speakers rather than memorized speech patterns, thus validating its ability to generalize beyond the training set.
This approach follows that used in [19] to ensure consistency with previous studies. The PVQD dataset originally contained 296 recordings, but 13 corrupted files were excluded, leaving a total of 283 valid samples. Specifically, 226 samples were used for training and 57 samples for testing. Table 2 provides the demographic distribution of speakers in the training and test sets, including gender ratio and age range. For a more comprehensive evaluation, vowel segments and connected-speech segments were extracted from each recording. Therefore, the PVQD dataset was divided into two subsets: PVQD-A (vowel-only subset), which contains /a/ vowel segments; and PVQD-S (speech-based subset), which consists of continuous speech segments. In this way, vowel phonation (PVQD-A) and connected-speech (PVQD-S) were evaluated independently. The number of samples in each subset is shown in Table 3. Since each recording contains one /a/ vowel and multiple continuous speech segments, there are more samples of continuous speech than /a/ vowel.
@@ 4.3 Feature Extraction
For the sustained-vowel subset (PVQD-A), recordings were made according to standard clinical instructions and were long enough to meet the requirement of $\geq$100 consecutive cycles for reliable perturbation analysis [6]. Perturbation measures were calculated based on stable voiced portions automatically identified by Parselmouth [43], excluding onset and offset frames.
Given the clinical limitations of perturbation analysis, it may not be applicable to running speech due to frequent voiced/unvoiced transitions and turbulent noise. For the connected-speech subset (PVQD-S) consisting of six standardized CAPE-V sentences, perturbation analysis was applied only to voiced frames to obtain auxiliary acoustic descriptors for modeling, rather than as clinical perturbation measures. Notably, the primary acoustic representation used by VOQANet+ is the SFM embedding, which captures the major phonatory and spectral cues. LLDs provide only supplementary information. Despite the limitations, combining these LLDs with SFM embeddings can provide complementary information and contribute to improved performance on sentence-based tasks.
@@ 4.6 Noise Robustness Setup
We introduced a noise-augmented version of the PVQD dataset to evaluate the robustness of the model under adverse acoustic conditions. Table 4 summarizes the noise types and signal-to-noise ratio (SNR) levels used for training, seen test, and unseen test scenarios. Each set contains the original clean utterances. The training set was augmented with four noise types: white, pink (colored), cafeteria babble, and cocktail party (background), at SNRs of –5, 0, 5, and 10 dB, covering both stationary and non-stationary interference commonly used in speech robustness research. To evaluate generalization, unseen noise types were added during testing: brown noise (low-frequency), baby cry, and laughter, the latter two reflecting real-world non-speech vocalizations relevant to clinical and telehealth settings, where voice assessment may be performed in varied background settings. Fig. 2 shows example waveforms and spectrograms for PVQD-A and PVQD-S under different noise conditions. The selected SNR range captures environments from challenging to moderately noisy, aligning with prior ASR and speech enhancement studies.
@@
Table 8: Cross-Validation Performance of VOQANet and VOQANet+.
Model Utterance-Level Patient-Level
PVQD-A PVQD-S PVQD-A PVQD-S
RMSE ↓ PCC ↑ RMSE ↓ PCC ↑ RMSE ↓ PCC ↑ RMSE ↓ PCC ↑
CAPE-V Prediction
[…]
GRBAS Prediction
[…]
@@
Table 10: Ablation Study on Different LLD Types Within The VOQANet+ Framework.
Feature PVQD-A PVQD-S
RMSE ↓ PCC ↑ RMSE ↓ PCC ↑
CAPE-V Prediction
CPP 13.194 0.750 12.804 0.765
JSH 11.555 0.787 14.864 0.708
WavLM(WS) + CPP 9.506 0.868 8.258 0.897
WavLM(WS) + JSH 9.042 0.878 7.356 0.908
GRBAS Prediction
CPP 0.423 0.714 0.418 0.717
JSH 0.398 0.772 0.518 0.685
WavLM(WS) + CPP 0.370 0.828 0.296 0.869
WavLM(WS) + JSH 0.332 0.839 0.289 0.874
Table 11: Generalization Performance of VOQANet and VOQANet+ on the Saarbrücken Voice Database (SVD).
Subset Accuracy F1-Score AUC
VOQANet (WavLM (WS))
[…]
VOQANet+ (WavLM (WS) + JSH)
[…]
@@ 5.1 Comparison of VOQANet with the Baseline
Table 5 presents the utterance-level performance of VOQANet using different SFM embeddings (last-layer HuBERT, WavLM, and Whisper) for CAPE-V and GRBAS prediction. Both HuBERT and WavLM are trained through SSL, whereas Whisper follows a semi-supervised training strategy. We compare VOQANet with the methods of Lin et al. [19], which evaluate traditional and neural regressors using LLDs, MFCC+MS, raw waveform features, and embeddings from pre-trained models such as W2V2, HuBERT, WavLM, and Whisper. Their evaluation focuses on CAPE-V prediction on the PVQD-A. Their results show that traditional features (HF and MFCC+MS) outperform most SFM features, with Whisper (Last) being the best SFM baseline (RMSE 14.76, PCC 0.69). However, their best results are significantly worse than those of VOQANet, suggesting that these traditional features are not sufficient to model the complex acoustic patterns relevant to pathological voice assessment. Notably, Lin et al. used noise-augmented training, whereas VOQANet in this experiment was trained only on clean PVQD-A data, yet still achieved far stronger results: RMSE to 12.921 (HuBERT (Last)), 10.514 (Whisper (Last)), and 9.955 (WavLM (Last)), with WavLM achieving the lowest RMSE and highest PCC (0.838).
On PVQD-S, VOQANet with WavLM (Last) also surpasses Whisper and HuBERT, suggesting that WavLM is particularly well suited to capture prosodic and phonetic nuances in continuous speech. Similar improvements were observed for GRBAS prediction on both subsets. Although GRBAS improvements appear smaller due to its narrower rating scale (0–3), they remain clinically meaningful.
@@ 5.2 Comparison of SFM Types and Representations
In this study, HuBERT, WavLM, and Whisper are used as backbone SFMs because of their excellent performance in speech quality assessment and the increasing relevance of SSL models in non-intrusive speech evaluation [38, 32]. To further analyze the effectiveness of SFM-based representations, we compare two adaptation strategies: last-layer features (Last) and weighted-sum aggregation (WS) for HuBERT, Whisper, and WavLM under utterance-level evaluation. As shown in Table 6, WS representation consistently outperforms the last-layer representation in all configurations. For instance, in GRBAS prediction on PVQD-A, WavLM (WS) yields lower RMSE (0.369 vs. 0.380) and higher PCC (0.809 vs. 0.795). This aligns with prior work [20, 21] showing that aggregating multi-layer transformer features is more effective than relying solely on the final layer. The last layer primarily captures high-level linguistic and semantic representations, while the intermediate layers retain lower-level acoustic and phonatory information such as spectral smoothness, perturbation, and harmonic balance, which are essential for pathological voice analysis. The WS aggregation effectively integrates these multi-level features, enabling the model to leverage both global and fine-grained information for more accurate prediction of perceptual voice quality. Similar advantages have been reported in recent audio quality studies [45]. This finding supports the motivation for adopting WS aggregation in subsequent experiments using WavLM embeddings.
@@
Table 6 also shows that WavLM outperforms HuBERT and Whisper on both PVQD-A and PVQD-S. For CAPE-V prediction on PVQD-S, WavLM (WS) achieves lower RMSE (9.209 vs. 9.933) and higher PCC (0.870 vs. 0.863) than Whisper, with even larger gains for GRBAS (PCC 0.845 vs. 0.822). These findings confirm that WavLM’s denoising pretraining and fine-grained acoustic modeling have advantages in disordered voice settings, especially when handling longer or more variable connected-speech contexts. Consequently, WavLM (WS) is adopted for subsequent experiments, which provides the most informative and powerful representation. Although prior studies identified HuBERT as particularly reliable for distinguishing normal from pathological voices [23], our findings show that Whisper and especially WavLM also deliver strong performance. This supports the clinical value of developing multiple reliable AI systems, as different SFMs may capture complementary acoustic cues relevant to dysphonia assessment. Therefore, in the following discussions, we only report WavLM as the representative for SSL model.